Find vertex and focus of parabola
4y = x squared - 12x + 52
How do you do this one?

Answer by stanbon(75887)

Find vertex and focus of parabola
4y = x squared - 12x + 52
----
Complete the square:
4y = x^2-12x+36 + 16
---
y = (1/4)(x-6)^2 + 4
---
Vertex: (6,4)
----
To find the focus:
4p = 1/4
p = 1/16
----
Focus is 1/16 above the vertex:
Focus: (6 , 4+(1/16)) or (6,67/16)
=========================================
Cheers,
Stan H.
=================